OpenStudy (anonymous):

Earth has a radius of about 6.4 x 10^6m. Approximate the surface area of Earth using the formula for the surface area of a sphere. S=4*pi*r^2

OpenStudy (anonymous):

S = 4 * π * (6.4 x 10^6m)^2 To square the radius remember that although it's scientific notation it's still really just multiplication. Remember the exponent rules for raising a multiplied quantity to a power: raise each thing being multiplied to that power S = 4 * π * 6.4^2 * (10^6m)^2 = 4 * 40.96 * 10^12 sq meters * π = 163.84 * 10^12 * π sq meters = 1.6384 * 10^14 * π sq meters approximately = 5.147 * 10^14 sq meters

OpenStudy (anonymous):



OpenStudy (anonymous):

oceans cover about 70% of the surface of the earth. about how many square meteres of earth's surface are covered by ocean water?

OpenStudy (anonymous):